Tammy found 1.95 in quarters and dimes in the car. There were four times as many dimes as quarters. How many dimes and quarters did she find?
Found 2 solutions by stanbon, shree840:
Answer by stanbon(57424)

Tammy found 1.95 in quarters and dimes in the car. There were four times as many dimes as quarters. How many dimes and quarters did she find?
--------------------
Equations:
Quantity: d = 4q
Value 10d + 25q = 195 cents
-------
Substitute and solve for "q":
10(4q) + 25q = 195
65q = 195
q = 3 (# of quarters)
--------
Solve for "d":
d = 4q
d = 4*3 = 12 (# of dimes)
============================
Cheers,
Stan H.
============

Answer by shree840(260)

4x*.1+x*.25=1.95
.65x=1.95
x=1.95/.65=3
so dimes 3*4=12=$1.20
quarters 3 =$.75